# LLM4Mat-Bench: Benchmarking Large Language Models for Materials Property Prediction

### v2

@@ E.2 Detailed MAE Results
Input Model Max CO2 $\downarrow$ Min CO2 $\downarrow$ LCD $\downarrow$ PLD $\downarrow$ Void fraction $\downarrow$ Surface area m2g $\downarrow$ Surface area m2cm3 $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 1.306 0.068 1.763 2.043 0.062 390.269 296.248 2.257
Composition Llama 2-7b-chat:0S 199.955 60.032 381.040 447.976 0.361 3239.754 2593.456 0.174
@@
Description Llama 2-7b-chat:0S 17.457 7.884 134.936 572.296 0.472 2956.530 2019.484 0.193
@@
Input Model FEPA $\downarrow$ EPA $\downarrow$ Ehull $\downarrow$ VPA $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 0.012 0.016 0.018 0.164 17.780
Composition Llama 2-7b-chat:0S 21.986 8.097 61.275 15.079 0.034
@@
Description Llama 2-7b-chat:0S 212.607 7.843 1.062 44.740 0.069
@@
Input Model Bandgap $\downarrow$ Total energy $\downarrow$ LCD $\downarrow$ PLD $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 0.351 261.601 0.965 0.795 3.076
Composition Llama 2-7b-chat:0S 0.947 1498.974 87.894 395.609 0.303
@@
Description Llama 2-7b-chat:0S 2.384 1791.591 154.422 539.926 0.152
@@
Input Model FEPA $\downarrow$ EPA $\downarrow$ Total energy $\downarrow$ Indirect Bandgap $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 0.642 3.155 7.883 0.038 61.729
Composition Llama 2-7b-chat:0S 438.919 1949.724 514.448 1.000 0.188
@@
Description Llama 2-7b-chat:0S 446.254 1763.805 182.817 1.316 0.264
@@
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ Wtd. Avg. (MAD/MAE) $\uparrow$
@@
Structure CGCNN (baseline) 0.033 0.062 14.496
Composition Llama 2-7b-chat:0S 38.194 2.166 0.105
@@
Description Llama 2-7b-chat:0S 29.677 2.221 0.106
@@
Input Model Bandgap $\downarrow$ MAD/MAE $\uparrow$
@@
Structure CGCNN (baseline) 0.292 2.751
Composition Llama 2-7b-chat:0S 0.907 0.885
@@
Description Llama 2-7b-chat:0S 5.194 0.155

### v3

@@ E.2 Detailed MAE Results
Input Model Max CO2 $\downarrow$ Min CO2 $\downarrow$ LCD $\downarrow$ PLD $\downarrow$ Void fraction $\downarrow$ Surface area m2g $\downarrow$ Surface area m2cm3 $\downarrow$ Wtd. Avg.
(MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 1.306 0.068 1.763 2.043 0.062 390.269 296.248 2.257
Comp. Llama 2-7b-chat:0S 199.955 60.032 381.040 447.976 0.361 3239.754 2593.456 0.174
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval. 34.676 2600.056 Inval. Inval.
Gemma 2-9b-it:5S 3.203 Inval. Inval. Inval. Inval. Inval. Inval. Inval.
@@
Descr. Llama 2-7b-chat:0S 17.457 7.884 134.936 572.296 0.472 2956.530 2019.484 0.193
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval. 12.033 2608.407 1649.961 Inval.
Gemma 2-9b-it:5S 3.203 Inval. Inval. Inval. Inval. Inval. Inval. Inval.
@@
Input Model FEPA $\downarrow$ EPA $\downarrow$ Ehull $\downarrow$ VPA $\downarrow$ Wtd. Avg. (MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 0.012 0.016 0.018 0.164 17.780
Comp. Llama 2-7b-chat:0S 21.986 8.097 61.275 15.079 0.034
@@
Descr. Llama 2-7b-chat:0S 212.607 7.843 1.062 44.740 0.069
@@
Mistral 7b-Instruct-v0.1:5S 0.866 11.713 0.463 26.56 0.147
Gemma 2-9b-it:5S 0.107 Inval. Inval. Inval. 1.013
@@
Input Model Bandgap $\downarrow$ Total energy $\downarrow$ LCD $\downarrow$ PLD $\downarrow$ Wtd. Avg. (MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 0.351 261.601 0.965 0.795 3.076
Comp. Llama 2-7b-chat:0S 0.947 1498.974 87.894 395.609 0.303
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval. Inval.
Gemma 2-9b-it:5S Inval. 688.079 Inval. Inval. Inval.
@@
Descr. Llama 2-7b-chat:0S 2.384 1791.591 154.422 539.926 0.152
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval. Inval. Inval. Inval.
Gemma 2-9b-it:5S Inval. 633.749 Inval. Inval. Inval.
@@
Input Model FEPA $\downarrow$ EPA $\downarrow$ Total energy $\downarrow$ Indirect Bandgap $\downarrow$ Wtd. Avg. (MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 0.642 3.155 7.883 0.038 61.729
Comp. Llama 2-7b-chat:0S 438.919 1949.724 514.448 1.000 0.188
@@
Descr. Llama 2-7b-chat:0S 446.254 1763.805 182.817 1.316 0.264
@@
Input Model FEPA $\downarrow$ Bandgap $\downarrow$ Wtd. Avg. (MAD:MAE) $\uparrow$
@@
CIF CGCNN (baseline) 0.033 0.062 14.496
Comp. Llama 2-7b-chat:0S 38.194 2.166 0.105
@@
Descr. Llama 2-7b-chat:0S 29.677 2.221 0.106
@@
Input Model Bandgap $\downarrow$ MAD:MAE $\uparrow$
@@
CIF CGCNN (baseline) 0.292 2.751
Comp. Llama 2-7b-chat:0S 0.907 0.885
@@
Mistral 7b-Instruct-v0.1:5S Inval. Inval.
Gemma 2-9b-it:5S Inval. Inval.
@@
Descr. Llama 2-7b-chat:0S 5.194 0.155